OpenStudy (anonymous):

Find vertex of x and y coordinate: f(x)=-2x^2+2x+3

OpenStudy (mertsj):

Do you mean find the x and y coordinates of the vertex?

OpenStudy (anonymous):

yes sorry

OpenStudy (mertsj):

\[x=\frac{-b}{2a}=\frac{-2}{4}=\frac{-1}{2}\]

OpenStudy (mertsj):

Now plug -1/2 into the equation and find f(x). You will have your vertex.

OpenStudy (mertsj):

\[2(\frac{-1}{2})^2+2(\frac{-1}{2})+3=\frac{1}{2}-1+3=\frac{5}{2}\]

OpenStudy (mertsj):

Vertex is (-1/2,5/2)

OpenStudy (anonymous):

Wouldn't it be 1/2 and not negative?

OpenStudy (mertsj):

Oh gees. Yes. I didn't see the negative sign in front of the 4. You'll have to recalculate. Sorry

OpenStudy (mertsj):

in front of the 2.

OpenStudy (anonymous):

No problem you have been a great help!

OpenStudy (mertsj):

ty